Suppose we are told that the speed, v, of a rotating particle depends on the radius, r,

and the frequency f, in the form v = 2πr

a

/fb

. If this is to be a dimensionally homogeneous

equation, determine the values of a and b.


1
Expert's answer
2020-08-19T14:07:21-0400

As per the given question,

Here the f is the frequency of the rotating particle and r is the radius and v is the velocity of the particle which is rotating in the circular path.

So, "v=\frac{2\pi ra}{fb}"

the dimensional formula of "v=[LT^{-1}]"

dimension formula of "r=[L]"

Dimensional formula of f is "[T^{-1}]"

Now, substituting the values,

"[LT^{-1}]=\frac{[L]a}{[T^{-1}]b}"

So, dimensional formula of a"=[T^{-1}]"

Dimension formula of b "=[T]"